# Optimized neural forms for solving ordinary differential equations

### v1

@@ Abstract
A critical issue in approximating solutions of ordinary differential equations using neural networks is the exact satisfaction of the boundary or initial conditions. For this purpose, neural forms have been introduced, i.e., functional expressions that depend on neural networks which, by design, satisfy the prescribed conditions exactly. Expanding upon prior progress, the present work contributes in three distinct aspects. First, it presents a novel formalism for crafting optimized neural forms. Second, it outlines a method for establishing an upper bound on the absolute deviation from the exact solution. Third, it introduces a technique for converting problems with Neumann or Robin conditions into equivalent problems with parametric Dirichlet conditions. The proposed optimized neural forms were numerically tested on a set of diverse problems, encompassing first-order and second-order ordinary differential equations, as well as first-order systems. Stiff and delay differential equations were also considered. The obtained solutions were compared against solutions obtained via Runge-Kutta methods and exact solutions wherever available. The reported results and analysis verify that in addition to the exact satisfaction of the boundary or initial conditions, optimized neural forms provide closed-form solutions of superior interpolation capability and controllable overall accuracy.
@@ 1 Introduction
Neural networks have been successfully employed for the solution of diverse problems in science and engineering. Their universal approximation property, when using activation functions of specific type [1, 2, 3], renders them particularly suitable for modeling tasks. Within this framework, neural networks have been employed to model solutions for both ordinary differential equations (ODEs) and partial differential equations (PDEs) as well. Although the first relevant neural methods appeared in the mid-90s [4, 5, 6, 7, 8, 9], their utilization remained limited for a period of about 20 years. This research stagnation can be attributed to several factors, including the lack of specialized and widely accessible software, as well as the limited computing power available at the time. Today, the rapid advances in deep learning, the emergence of powerful software platforms (TensorFlow, Keras, PyTorch, etc.) and the advent of high-performance computing technologies (GPUs, parallel multicore multiprocessor clusters), have brought neural differential equation solvers back into play. Thus, a number of alternative techniques, based on a plethora of neural network types, have been developed and are currently available [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], along with new finite difference based solvers such as ODIL [21].
@@
The primary goal of neural methods is the construction of a parametric model that can be tuned to produce a solution with a prescribed degree of accuracy. This is carried out by minimizing a properly designed error function (also known as loss function or cost function) within the space of the model parameters. In addition, differential equations are accompanied by boundary or initial conditions that the solution must satisfy. In the following, we will use the term “the conditions” to refer to either the boundary conditions or the initial conditions, unless otherwise stated. It is important to realize that a trial solution which fails to meet the prescribed conditions, cannot be considered as an adequate approximation to the true solution of the problem under consideration.
@@
A common approach treats the conditions as equality constraints when minimizing the error function. In this case, constrained optimization methods (interior point, active set, augmented Lagrangian, etc.) that exactly satisfy the constraints should be used. However, their application appears to be complicated and time-consuming as well. Alternatively, equality constraints may be approximately treated via penalty methods using unconstrained optimization. Penalty terms are added to the original error function in an attempt to inhibit constraint violations. In fact, this is by far the most commonly followed approach, probably due to reasons of simplicity and implementation convenience. If the parametric trial solution is crafted in such a way that it exactly satisfies the specified conditions, regardless of the parameter values, the constraints are eliminated, thereby allowing for the application of more efficient unconstrained optimization methods. In the case of ODEs, as well as of PDEs defined inside rectangular domains, this approach is preferable because the exact satisfaction of the conditions allows the method to attain higher accuracy as pointed out in [22]. In contrast, for PDEs within complex geometry boundaries, the design of proper trial solutions is quite challenging, thereby rendering the approximate penalty approach rather appealing.
@@
In cases where the conditions are treated as constraints, the trial solution typically consists of a single neural network. On the other hand, the constraint elimination approach requires more elaborate functional constructs. To this end, neural forms were indirectly introduced in [8] and further refined in [17]. In both cases, a rigid polynomial expression was used to account for the conditions. An alternative neural form type was introduced in [10], involving a function calculating the distance of a domain point to the boundary, an idea that was further extended in [14] for the case of Dirichlet conditions.
Obviously, neural forms that exactly satisfy the ODE conditions are not unique. Therefore, the question concerning the existence of an optimal (in a certain sense) choice arises spontaneously. This is one of the issues addressed here, by introducing the “optimized neural forms”. More specifically, the contributions of the present work are as follows:
@@
A procedure is introduced for obtaining an estimate of the quality of the approximate solution, namely a reliable upper bound of its absolute deviation from the exact solution.
@@
A technique is devised transforming Neumann and Robin conditions into equivalent Dirichlet parametric conditions.
The performance of the proposed optimized neural forms was assessed on first-order and second-order ODEs, as well as on first-order systems, under Dirichlet, Neumann, mixed Dirichlet-Neumann, Robin, and Cauchy conditions. The obtained solutions were compared to exact solutions (wherever available), to solutions obtained by the penalized neural method, as well as to solutions obtained by the fourth-order Runge-Kutta method. The relevant analysis verified that the proposed optimized neural forms can yield superior closed-form solutions, while exactly satisfying the prescribed conditions.
The rest of the article is organized as follows: Section 2 briefly reviews neural ODE solvers, and introduces the neural forms concept. Section 3 elaborates on the optimized neural forms framework for first-order and second-order ODEs. Section 4 analyzes the transformation of Neumann/Robin conditions into equivalent Dirichlet parametric conditions. Section 5 derives an upper bound of the absolute deviation between the approximate and the exact solution. Section 6 describes in detail the experimental setting. Section 7 presents the obtained solutions, and compares them to exact solutions as well as to solutions obtained by the Runge-Kutta method. Section 8 summarizes the contributions of this work, draws some conclusions, and suggests further research.
@@ 2 Neural forms for solving ordinary differential equations
Let the general form of a second-order ODE:
with boundary conditions:
or initial conditions:
@@
This problem is typically addressed by minimizing the left-hand side of Eq. (5) subject to the conditions of Eq. (6) or (7) using constrained optimization methods. In case the baseline neural method with penalties is employed, as in [20, 16], the error function takes the following form:
where is the penalty coefficient.
Alternatively, neural forms assume that the trial solution may be written as the sum of two terms:
where is any smooth function that matches the boundary or initial conditions, henceforth referred to as boundary match or initial match, respectively. Moreover, is a parametric function, which typically involves a neural network, with vanishing contribution to the respective conditions regardless of the value of its parameters , i.e.:
Note that, under the above considerations, the trial solution satisfies the ODE conditions by construction.
The functions and are closely related and their design is based on a common operator. Let be a boundary match operator defined as:
This operator is not unique. In fact, there is an infinite number of alternative forms. For instance, for the case of Dirichlet boundary conditions, it may be given by a linear relation as:
@@
with respect to the weights .
3 Optimized neural forms for different types of boundary and initial conditions
As illustrated in Eqs. (10) and (11), the function is not unique. Therefore, we can take a further step by assuming that it is a parametric function , satisfying the problem’s conditions for any values of the parameters . The error function of Eq. (12) can then be rewritten as:
which is now minimized with respect to both and , leading to an optimized choice for the match and, therefore, for the neural form solution as well. In the following paragraphs, we elaborate on this concept for the cases of first-order ODEs and systems, as well as for second-order ODEs.
3.1 Optimized neural forms for first-order ordinary differential equations and systems
@@
with the initial condition . Then the functions , , and , of the neural form can be given as:
where is a bounded function in . Note that may be a neural network, i.e., . The associated error function is given as follows:
@@
Then, the initial matches are defined as:
@@
3.2 Optimized neural forms for second-order ordinary differential equations
Second-order ODEs may be accompanied by Dirichlet, Neumann, mixed Dirichlet-Neumann, Robin, or Cauchy conditions. Each condition type corresponds to a different boundary match operator, hence, we examine each case separately.
3.2.1 Dirichlet conditions
@@
where is a second-order differential operator. Also, consider a function defined for all , with:
There are diverse ways to construct . For example, it may be given as:
with and being functions in . Let us employ the form of Eq. (16) with:
@@
which is the derived form for the Dirichlet conditions.
3.2.2 Mixed Dirichlet-Neumann conditions
3.2.3 Neumann conditions
3.2.4 Cauchy conditions
3.2.5 Robin conditions
4 Reductive transformation of Neumann to equivalent Dirichlet boundary conditions
Instead of maintaining different boundary match operators for all condition types, neural forms for Dirichlet conditions, i.e., and , can be solely considered. In this case, a mixed problem with conditions and can be tackled by replacing with a suitable expression, such that the Dirichlet-type neural form satisfies the mixed conditions.
Let us demonstrate this idea with a simple example using the following non-parametric boundary match operator:
@@ 4 Reductive transformation of Neumann to equivalent Dirichlet boundary conditions
Demanding that and solving for gives:
Substituting in Eq. (23) results in the neural form that satisfies the mixed boundary conditions.
We provide the corresponding result for the case of Robin conditions of Eq. (21) for the non-parametric boundary match operator of Eq. (22):
By choosing appropriate values for , , , and , in the above formulation, we can produce the corresponding formulae for all combinations of boundary conditions.
5 Solution quality assessment
Assessing solution quality is a challenging issue when solving ODEs numerically. We present a perturbation approach that can yield, under certain weak assumptions, a reliable upper bound for the absolute deviation between the obtained and the actual solution:
More specifically, let:
be the trial solution obtained by an optimized neural form. Assuming that the error function in eq. (13) is approximated by the mean squared error, the corresponding error value is positive and can be written as:
where is a scalar. We construct a new trial solution by adding a perturbation to the obtained trial solution . In order to preserve condition satisfaction, the perturbation should be of the form:
with being a neural network. The perturbed trial solution corresponds to an error function that is minimized with respect to the perturbation parameters , starting from an initial vector with the property:
This way, we arrive at a lower error value, , corresponding to an optimum vector .
If the studied ODE has the form:
then the perturbed error value is:
Assuming that is a small perturbation, we consider the first-order approximation:
@@ 5 Solution quality assessment
which is a reliable estimate for the upper bound , as long as is a small perturbation.
@@ 6 Experimental setting
In the following paragraphs, details regarding the experimental setting for the assessment of the proposed optimized neural forms are provided. More specifically, information regarding the employed test problems, dataset generation, neural network architecture, and solution quality criteria are given.
6.1 Test suite
The test suite considered for the experimental evaluation of the optimized neural forms embraced a number of first-order and second-order ODEs, as well as first-order ODE systems, equipped with diverse types of boundary and initial conditions. Details regarding the employed test problems are provided in the following paragraphs.
@@ 6.1 Test suite
Test Problem 2 (second-order ODE)
This is a nonlinear ODE defined as:with exact solution:
It has been considered as a boundary value problem with Dirichlet, Neumann, mixed Dirichlet-Neumann, and Robin conditions, as well as an initial value problem with Cauchy conditions.
@@
This problem is derived by Bessel functions of first kind , which are defined as solutions of the Bessel ODE subject to Cauchy initial conditions:(28) which are nonsingular at the origin. Here, the case was considered for .
@@
This is a stiff system defined as [23]:(30) and exact solution:
@@
This is a nonlinear system derived by reducing the second-order ODE of Test Problem 2:(31) This problem was considered under various types of boundary conditions, namely Dirichlet, Neumann, mixed Dirichlet-Neumann, as well as with Cauchy initial conditions.
7.
Problem 7 (single-delay ODE)
This is a single-delay ODE defined as [24]:(32) and exact solution:
In this case, the trial solution is the same as that for a first-order ODE without delays, i.e.:
However, in order to account for the delay term, the error function becomes:
The exact solutions for all test problems are illustrated in Fig. 1.
| Training points () | Neural parameters | ||||||
| Problem | low | medium | high | Test points () | low | medium | high |

### v2

# Augmented neural forms with parametric boundary-matching operators for solving ordinary differential equations

@@ Abstract
Approximating solutions of ordinary and partial differential equations constitutes a significant challenge. Complementarily to numerical analysis methods, neural forms have emerged as a valuable solution methodology to this problem. Based on functional expressions that inherently depend on neural networks, neural forms are specifically designed to precisely satisfy the prescribed initial or boundary conditions of the problem, while providing the approximate solutions in closed form. Departing from the fundamentally important class of ordinary differential equations, the present work aims to refine and validate the neural forms methodology, paving the ground for further developments in more challenging fields. The main contributions are as follows. First, it introduces a formalism for systematically crafting proper neural forms with adaptable boundary matches that are amenable to optimization. Second, it describes a novel technique for converting problems with Neumann or Robin conditions into equivalent problems with parametric Dirichlet conditions. Third, it outlines a method for determining an upper bound on the absolute deviation from the exact solution. The proposed augmented neural forms approach was tested on a set of diverse problems, encompassing first- and second-order ordinary differential equations, as well as first-order systems. Stiff differential equations have been considered as well. The resulting solutions were subjected to assessment against existing exact solutions, solutions derived through the common penalized neural method, and solutions obtained via contemporary numerical analysis methods. The reported results demonstrate that the augmented neural forms not only satisfy the boundary and initial conditions exactly, but also provide closed-form solutions that facilitate high-quality interpolation and controllable overall precision. These attributes are essential for expanding the application field of neural forms to more challenging problems that are described by partial differential equations.
@@ 1 Introduction
Neural networks have been successfully employed for the solution of diverse problems in science and engineering. Their universal approximation property, when using activation functions of specific type [1, 2, 3], renders them particularly suitable for modeling tasks. Within this framework, neural networks have been employed to model solutions for ordinary differential equations (ODEs) and partial differential equations (PDEs) as well. Although the first relevant neural methods appeared in the mid-90s [4, 5, 6, 7, 8, 9], their utilization remained limited for a period of about 20 years. This stagnation can be attributed to several factors, including the lack of widely accessible specialized software, as well as the limited computing power available at the time. The recent advancements in deep learning, the emergence of sophisticated software platforms (such as TensorFlow, Keras, and PyTorch) and the advent of high-performance computing technologies (including GPUs and parallel multicore multiprocessor clusters) have led to a resurgence of interest in neural differential equation solvers. Thus, a number of alternative techniques, based on a plethora of neural network types, have been developed and are currently available [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], along with new finite difference based solvers such as ODIL [21].
@@
The primary goal of neural methods is the construction of a parametric model that can be tuned to produce a solution with a prescribed degree of accuracy. This is carried out by minimizing a properly designed error function (also known as loss or cost function) within the space of the model parameters. In addition, differential equations are accompanied by boundary and/or initial conditions that must be satisfied. From here on we will use the term conditions to refer either to boundary or initial conditions, unless otherwise specified. It is important to realize that a trial solution which fails to meet the prescribed conditions cannot be considered as an adequate approximation to the true solution of the problem under consideration.
@@
A common approach treats the conditions as equality constraints when minimizing the error function. In this case, constrained optimization methods (interior point, active set, augmented Lagrangian, etc.) that exactly satisfy the constraints should be used. However, their application appears to be complicated and time-consuming as well. Alternatively, equality constraints may be approximately treated via penalty methods using unconstrained optimization. Penalty terms are added to the original error function in an attempt to inhibit constraint violations. In fact, this is by far the most commonly followed approach, perhaps due to reasons of implementation convenience and simplicity of interpretation. If the parametric trial solution is crafted in such a way so as to satisfy exactly the specified conditions regardless of the parameter values, the constraints are eliminated, thereby allowing for the application of efficient unconstrained optimization methods. In the case of ODEs, as well as of PDEs that are defined inside rectangular domains, this approach is preferable because the exact satisfaction of the conditions allows the method to attain higher accuracy as pointed out in [22]. In contrast, for PDEs within domains of complex geometry, the design of proper trial solutions that satisfy prescribed conditions is challenging, and thus the alternative approximate penalty approach has been frequently preferred.
@@
In cases where the conditions are treated as constraints, the trial solution typically consists of a single neural network. On the other hand, the constraint elimination approach requires more elaborate functional constructs. To this end, neural forms were indirectly introduced in [8] and further refined in [17]. In both cases, a rigid polynomial expression was used to account for the conditions. An alternative neural form type was introduced in [10], involving a function representing the distance of a domain point from the boundary, an idea that was further extended in [14] for the case of Dirichlet conditions. Neural forms that exactly satisfy the ODE conditions are obviously not unique. Therefore, the question concerning the existence of a possibly optimal choice naturally arises. This is one of the issues addressed here by introducing the augmented neural forms that are presented in the following paragraphs.
The contributions of the present work can be summarized as follows:
@@
A technique is introduced for transforming Neumann or Robin to equivalent parametric Dirichlet conditions.
@@
A methodology is introduced for evaluating the quality of the approximate solution by determining a reliable upper bound of its absolute deviation from the exact solution.
The performance of the proposed approach was assessed on first- and second-order ODEs, as well as on first-order systems, subject to Dirichlet, Neumann, mixed Dirichlet-Neumann, Robin, and Cauchy conditions. The obtained solutions were compared to existing exact solutions, to solutions obtained by the penalized neural method, and to solutions obtained by the ode113 numerical solver of Matlab® [23], which is based on a state-of-the-art variable-step, variable-order Adams-Bashforth-Moulton method. The conducted analysis verified that the proposed augmented neural forms can provide robust, closed-form solutions with exact satisfaction of the prescribed conditions, along with excellent generalization performance.
The rest of the article is organized as follows: Section 2 briefly reviews neural ODE solvers, and introduces the neural forms concept. Section 3 elaborates on the augmented neural forms framework for first- and second-order ODEs. Section 4 analyzes the transformation of Neumann or Robin conditions to equivalent parametric Dirichlet conditions. Section 5 provides a comprehensive analysis of obtaining a reliable upper bound for the absolute deviation between the approximate and the exact solution. Section 6 describes in detail the experimental setup. Section 7 presents the obtained solutions and a comparison thereof, as mentioned above. Finally, Section 8 provides an overview of the contributions of this work, along with comments, conclusions, and suggestions for further research.
@@ 2 Neural forms for solving ordinary differential equations
Consider the general form of a second-order ODE:
with boundary conditions (BCs):
or initial conditions (ICs):
@@
This problem is typically addressed by minimizing the left-hand side of Eq. (5) under the conditions of Eqs. (6) or (7), using constrained optimization methods. The baseline neural method [20, 16], adds penalties to the error function which assumes the following form:
and employs unconstrained optimization, with being a positive penalty-regulating coefficient.
Alternatively, neural forms cast the trial solution as a sum of two terms:
where is a smooth function matching the boundary or initial conditions, henceforth referred to as the boundary match or initial match, respectively. Moreover, is a parametric function typically depending on a neural network with null contribution to the problem’s conditions, i.e.:
Note that, with the above formulation, the trial solution satisfies the prescribed ODE conditions by construction.
The functions and are closely related, and their design is based on a common matching operator. Let be a boundary matching operator defined as:
We can easily verify that this operator is not unique. In fact, there is an infinite number of alternative forms. For instance, for the case of Dirichlet boundary conditions, may be given by a linear relation as follows:
@@
with respect to the neural network’s weight vector .
3 Augmented neural forms for different types of boundary and initial conditions
As illustrated in Eqs. (10) and (11), the function is not unique. Therefore, we can take a further step by representing it as a parametric function that satisfies the problem’s conditions for any values of the parameters vector . The error function of Eq. (12) can then be written as:
and minimized with respect to both and , leading to an optimized choice for the match and, consequently, for the trial neural form solution. In the following paragraphs, we elaborate on this concept for the cases of first-order ODEs and systems, as well as for second-order ODEs.
3.1 Augmented neural forms for first-order ordinary differential equations and systems
@@
with the initial condition . The functions , , and , are then given by:
where is a smooth, bounded function for . Note that may be a neural network, i.e., . The associated error function is then given by:
@@
with , where:
and . Then, the initial matches are defined as:
@@
3.2 Augmented neural forms for second-order ordinary differential equations
Second-order ODEs may be accompanied by Dirichlet, Neumann, mixed Dirichlet-Neumann, Robin, or Cauchy conditions. Each condition type corresponds to a different boundary matching operator, hence, we examine each case separately.
3.2.1 Case of Dirichlet conditions
@@
where is a second-order differential operator. Also, consider a function defined for all , satisfying and Then, there are diverse ways to construct . For example, it may be crafted as:
with and be smooth, bounded functions in . Let us employ the form of Eq. (16) with:
@@
which is a proper neural form for the case of Dirichlet conditions.
3.2.2 Case of mixed Dirichlet-Neumann conditions
3.2.3 Case of Neumann conditions
3.2.4 Case of Cauchy conditions
3.2.5 Case of Robin conditions
4 Reductive transformation of Neumann conditions to equivalent Dirichlet boundary conditions
Instead of maintaining different boundary match operators for all condition types, neural forms for Dirichlet conditions of the form , , can be solely considered. In this case, a mixed problem with conditions and can be tackled by replacing with a suitable expression, such that the Dirichlet-type neural form satisfies the mixed conditions. Let us demonstrate this idea with a simple example, using the following non-parametric (rigid) boundary matching operator:
@@ 4 Reductive transformation of Neumann conditions to equivalent Dirichlet boundary conditions
Demanding that and solving for yields:
Substituting in Eq. (23) results in the neural form that satisfies the mixed boundary conditions. We provide the following result:
for the case of Robin conditions of Eq. (21).
5 Goodness of the approximate solution
Assessing the quality of an approximate solution is of fundamental importance. Relevant research outcomes have recently been reported in [24]. In the following paragraphs, we present a perturbation approach that can yield, under certain weak assumptions, a reliable upper bound for the absolute deviation of the obtained solution from the exact one. More specifically, let:
be the trial solution obtained by an augmented neural form. The corresponding error is positive and may be written as , where is a scalar. We now construct a new trial solution by adding a perturbation to the obtained trial solution . In order to preserve the satisfaction of the prescribed conditions, the perturbation should be of the form:
with be a neural network. The perturbed trial solution, i.e., , produces the error function that is minimized with respect to the perturbation parameters , starting from an initial vector that satisfies . The minimized error value, corresponding to an optimum vector , is .
Consider a second order ODE of the form . The perturbed error value is then given as:
Assuming that is a small perturbation, we may consider the first-order approximation:
@@ 5 Goodness of the approximate solution
This provides a reliable estimate for an upper bound of as long as the assumed linear approximation in Eq. (25) is valid. In turn, this implies that must be a small perturbative correction.
@@ 6 Experimental setting
The following section provides a comprehensive presentation of the experimental assessment procedure of the proposed augmented neural forms. More specifically, it includes detailed information on the selected test problems, the dataset generation schemes, the neural architecture, and the performance measures.
6.1 Test problems
The test problems included a number of first-order and second-order ODEs, as well as first-order ODE systems, accompanied by boundary and initial conditions of various types that are frequently encountered in real world problems.
@@ 6.1 Test problems
Test Problem 2 (second-order, nonlinear ODE)
This problem is defined as:with exact solution:
In our experiments, it has been considered as a boundary value problem with Dirichlet, Neumann, mixed, and Robin conditions, and also as an initial value problem with Cauchy conditions.
(a) Test Problem 1
(b) Test Problem 2
(c) Test Problem 3
(d) Test Problem 4
(e) Test Problem 5
(f) Test Problem 6 Figure 1: Plots of the exact solutions for all test problems. -
@@
This is Bessel’s equation subject to Cauchy initial conditions:(29) The exact solution is Bessel’s function of first kind , which is non-singular at the origin. Here, the case was considered.
@@
This is a stiff system defined as [25]:(31) with exact solution:
@@
This is a nonlinear system derived by reducing the second-order ODE of Test Problem 2 as follows:(32) This problem was considered under various types of boundary conditions, namely Dirichlet, Neumann, mixed Dirichlet-Neumann, as well as with Cauchy initial conditions.
The exact solutions for all test problems are depicted in Fig. 1.
6.2 Dataset generation
For each test problem, two sets of data points were generated, namely a training dataset of size that was used to train the neural forms, and a test dataset of size that was used to assess the generalization quality. For the training dataset generation, two grid types were considered with data points :
1.
Equidistant points:
2.
Chebyshev points:
The two grid types offer different point distributions inside the interval of interest. Chebyshev grids get denser towards the end points and are known to suppress the Runge phenomenon. Gauss-Legendre grids were tested as well, offering almost identical results to those of the Chebyshev grid. Hence, they were omitted from our study.
| Test | Training points () | Test points () | Neural parameters () | ||||
| Problem | low | medium | high | low | medium | high | |